a hand grenade explodes into 3 pieces each piece flying 120 degrees relative to each other.

Piece 1 has a mass of 4 kg --

Piece 2 has a mass of 8 kg & speed of 2 m/s

Piece 3 has a mass of 8 kg & speed of 2 m/s

what is the speed of piece 1?
Answer #1

Given that after the explosion, each piece flies away at 120 degrees relative to each other.

Let's assume that the 4kg piece moves along the x-axis.

So the other 2 pieces fly away at 120 deg and -120deg(240 deg) respectively.

==============

Consider piece 2

v2sin0j = 2m/s* cos120i 2m/s sin1201 2 2r V2y= vcosi

v2=11 +1.732

---------------------------

Consider piece 3

3yJ= v3cosi +v3sin0j = 2m/s* cos240i2m/s* sin240j 33r

v3=-1i- 1.732

===========

During the explosion momentum is conserved

Initial momentum = Final momentum

Mum1 +m2u2 +m3U3

4kg 18kg(-li1.732j) 8kg* (-li- 1.732j) 20* Om/s

0=4*v_{1}+8*(-1i+1.732j)+8*(-1i-1.732j)

4*v_{1}+(-8i+13.856j)+(-8i-13.856j)=0

4*v_{1}+-8i+13.856j+-8i-13.856j=0

4*v_{1}+(-8-8)i+(13.856-13.856)j=0

4*v_{1}+-16i+0j=0

4*v_{1}=16i+0j

v_{1}=\frac{16i+0j}{4}

v_{1}=4i+0j

ANSWER: {\color{Red} v_{1}=4m/s}

=====================